# Short-video Propagation Influence Rating: A New Real-world Dataset and A New Large Graph Model

### v1

@@
Abstract.
Short-video platforms have gained immense popularity, captivating the interest of millions, if not billions, of users globally. Recently, researchers have highlighted the significance of analyzing the propagation of short-videos, which typically involves discovering commercial values, public opinions, user behaviors, etc. This paper proposes a new Short-video Propagation Influence Rating (SPIR) task and aims to promote SPIR from both the dataset and method perspectives. First, we propose a new Cross-platform Short-Video (XS-Video) dataset, which aims to provide a large-scale and real-world short-video propagation network across various platforms to facilitate the research on short-video propagation. Our XS-Video dataset includes 117,720 videos, 381,926 samples, and 535 topics across 5 biggest Chinese platforms, annotated with the propagation influence from level 0 to 9. To the best of our knowledge, this is the first large-scale short-video dataset that contains cross-platform data or provides all of the views, likes, shares, collects, fans, comments, and comment content. Second, we propose a Large Graph Model (LGM) named NetGPT, based on a novel three-stage training mechanism, to bridge heterogeneous graph-structured data with the powerful reasoning ability and knowledge of Large Language Models (LLMs). Our NetGPT can comprehend and analyze the short-video propagation graph, enabling it to predict the long-term propagation influence of short-videos. Comprehensive experimental results evaluated by both classification and regression metrics on our XS-Video dataset indicate the superiority of our method for SPIR.
1. Introduction
Short-video platforms, such as Tiktok (aka Douyin)111https://www.tiktok.com/, Kwai (aka Kuaishou)222https://www.kwai.com/, and Xigua333https://www.ixigua.com/, have gained immense popularity, captivating the interest of millions, if not billions, of users globally. The users are capable of publishing their activities, opinions, and thoughts by posting, commenting, and sharing various forms of online content, including text, images, and videos. Consequently, these overwhelming user behaviors and user-generated content have formed a huge and intricate propagation network, harboring a wealth of information about our society. Recently, researchers have highlighted the significance of analyzing the propagation of short-videos (Tang et al., 2022; Zhong et al., 2024; Song and Liu, 2024), which typically involves discovering commercial values, public opinions, user behaviors, etc. Traditional research mainly focuses on forecasting the popularity of videos (Chen et al., 2016; Xie et al., 2020, 2021) measured by a single indicator, such as the number of views or likes. However, real-world propagation behaviors are more complex and can include sharing, collecting, commenting, etc.
@@ 1. Introduction
Daum (Cha et al., 2009) | 1 | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | 196,037 | 196,037 | Unknown |
Youtube (Borghol et al., 2012) | 1 | ✓ | ✓ | ✗ | ✗ | ✓ | ✓ | ✗ | 1,761 | 1,761 | Unknown |
NUS (Chen et al., 2016) | 1 | ✓ | ✓ | ✗ | ✗ | ✓ | ✓ | ✗ | 303,242 | 303,242 | 98,166 |
Xigua (Xie et al., 2020) | 1 | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | 11,219 | 11,219 | 2,664 |
MicroLens (Zhong et al., 2024) | 1 | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | 19,738 | 19,738 | 100,000 |
@@
To address the above limitation, we attempt to evaluate the propagation influence by simultaneously considering as many interactive indicators as possible, including views, likes, shares, collects, comments, etc. Moreover, as shown in Figure 1, we propose a new task named Short-video Propagation Influence Rating (SPIR), which aims to predict the long-term propagation influence of a newly posted short-video, under the background of a huge propagation network. Different from traditional popularity prediction that forecasts a single interactive indicator over a short period (e.g., several hours (Xie et al., 2021; Tang et al., 2022)), our SPIR predicts the comprehensive propagation influence assessed through multi-dimensional interactions accumulated over a long period. SPIR requires a model to thoroughly analyze the multimodal content features, interactive information, and propagation relationships within a short-video. Subsequently, a rating from 0 to 9 is assigned to represent the propagation influence of this video that can be achieved over a long period. The development of SPIR systems can potentially benefit a wide spectrum of real-world applications on short-video platforms, including advertising (Yuan et al., 2022; Xiao et al., 2023), recommendation (Cai et al., 2023; Zhang et al., 2024a), and democracy (Chen et al., 2021; Qin et al., 2023). In this paper, we attempt to promote the research of SPIR from both the data and model perspectives, proposing the first cross-modal short-video propagation dataset and a novel large graph model for SPIR.
@@
To promote the research of SPIR and short-video propagation analysis, we propose a large-scale Cross-platform Short-Video (XS-Video) dataset, where videos are collected from the 5 biggest Chinese platforms (i.e., Douyin, Kuaishou, Xigua, Toutiao, and Bilibili). Though hot topics and content are typically discussed and replicated across platforms, almost all existing datasets collect data from a single platform and lack a broad scope of short-videos. As compared in Table 1, our XS-Video dataset significantly surpasses the existing datasets in terms of platform numbers and interactive indicator coverage, while also containing comment content and more samples and users. Specifically, we collect 381,926 states/samples of 117,720 short-videos within several days after their publication, belonging to 535 topics. Different states/samples of the same video are distinguished by their sampling times, as well as the interactions and comment content at those times. Furthermore, we annotate the propagation influence level of all short-videos according to their multi-dimensional interactive indicators over a two-week period following their publication, since almost all short-videos peak in interactions within two weeks. We demonstrate a short-video sample in the XS-Video dataset in Figure 2. Moreover, utilizing the entities and relations present in XS-Video, we construct a huge propagation graph consisting of 5.5 million nodes of videos, topics, comments, various interactions, etc., and 1.7 billion directed edges between these nodes. More details about our dataset and the constructed propagation graph are provided in Appendix A. Compared to existing datasets, XS-Video provides more comprehensive information on video content, post, interactions, and comments, which is essential for the research on short-video propagation, including our SPIR task.
@@
We build a comprehensive benchmark for SPIR by evaluating state-of-the-art methods, including Graph Neural Networks (GNNs), Large Language Models (LLMs), and multimodal LLMs. Surprisingly, these state-of-the-art methods are far from satisfactory for short-video propagation analysis. To improve the accuracy of SPIR and leverage the advancements in LLMs, we propose a Large Graph Model (LGM) named NetGPT. Our model can capture both the propagation network structure and heterogenous short-video features while utilizing the pretraining knowledge to predict the propagation influence level. LLMs have recently shown powerful reasoning ability for text, images, and even videos (Liu et al., 2024; Wu et al., 2024; Xue et al., 2024; Wang et al., 2024a; Li et al., 2024b). Current research has explored adapting LLMs for analyzing textual graph data (Tang et al., 2024; Li et al., 2024a; Wang et al., 2024b). However, bridging heterogeneous graph data and LLMs is still an open problem, especially as we cannot input our huge propagation graph into an LLM. Our experiments also reveal that LLMs cannot directly learn the graph-structured relations in short-video propagation data, which results in poor performance for SPIR. To address this limitation, we propose a three-stage training mechanism for establishing an effective LGM, including Heterogeneous Graph Pretraining, Supervised Language Fine-tuning, and Task-oriented Predictor Fine-tuning. Finally, we can empower LLMs (e.g., Qwen2-VL (Wang et al., 2024a) in our experiments) to comprehend and analyze the short-video propagation graph. Our NetGPT model significantly outperforms state-of-the-art baselines, including Graph Neural Networks (GNNs), LLMs, and multimodal LLMs on our XS-Video dataset. Our dataset and code are available at https://github.com/LivXue/short-video-influence.
@@
We propose the Short-video Propagation Influence Rating (SPIR) task, which aims to predict the long-term propagation influence of a newly posted short-video in a huge propagation network. Different from traditional popularity prediction that forecasts a single interactive indicator over a short period (e.g., several hours (Xie et al., 2021; Tang et al., 2022)), SPIR predicts the comprehensive propagation influence assessed through multi-dimensional interactions over a long period.
@@
2. Related Work
2.1. Short-Video Propagation Influence Analysis
Existing research on short-video propagation influence analysis mainly focuses on the popularity prediction of short-videos, such as predicting the views of newly posted videos (Roy et al., 2013; Chen, 2016; Jing et al., 2017; Tang et al., 2017; Ou et al., 2022). For example, Chen et al. (Chen et al., 2016) focus on the short-video representation and present a transductive multi-modal learning model for short-video popularity prediction. Xie et al. (Xie et al., 2020) propose a multimodal variational encoder-decoder framework that considers the uncertain factors as the randomness for the mapping from the multimodal features to the popularity. Xie et al. (Xie et al., 2021) propose a hierarchical multimodal variational encoder-decoder to predict the popularity of short-videos by comprehensively leveraging user information and the short-video content in a hierarchical fashion. Zhou et al. (Zhong et al., 2024) present a multimodal retrieval-augmented popularity prediction model that improves prediction accuracy using relevant retrieved information.
@@ 2.1. Short-Video Propagation Influence Analysis
2.2. Large Graph Model
Recent advancements in Large Language Models (LLMs) have showcased their impressive reasoning abilities not only with text but also with images and videos (Hu et al., 2024b; Ge et al., 2024; Cha et al., 2024; Li et al., 2025). This progress has led to an increasing focus on adapting LLMs for analyzing graph-structured data, especially textual graphs. For instance, Li et al. (Li et al., 2024a) utilize LLM prompting on textual graphs, offering substantial potential to enhance graph transfer capabilities across diverse tasks and domains. Tang et al. (Tang et al., 2024) integrate LLMs with graph structural knowledge through graph instruction tuning, which includes a text-graph grounding component, a dual-stage instruction tuning approach, and a graph-text alignment projector. Wang et al. (Wang et al., 2024b) propose Build-a-Graph Prompting and Algorithmic Prompting, two instruction-based approaches to enhance LLMs in solving natural language graph problems.
@@ 2.2. Large Graph Model
3. Dataset
Analyzing propagation patterns on short-video platforms can aid in discovering high-value content, public opinions, and user preferences, thereby creating commercial and social value. Our proposed dataset aims to provide a realistic and large-scale benchmark for the research on short-video propagation. As briefly shown in Table 1 and Figure 2, compared to previous short-video propagation datasets (Cha et al., 2009; Borghol et al., 2012; Chen et al., 2016; Xie et al., 2020; Zhong et al., 2024), our dataset has three key advantages: (1) While existing datasets separately collect data from a single short-video platform, we collect short-videos from the 5 biggest Chinese platforms, which provide cross-platform propagation information; (2) Instead of focusing on a subset of interactive indicators, we gather a wide range of indicators such as views, likes, shares, collects, fans, and comments, which facilitate deep analysis of video interactions; (3) Our dataset contains more samples and users, while providing more comprehensive short-video content, including videos, titles, descriptions, topics, and the related comment content. Figure 3 demonstrates a brief construction procedure of our dataset.
Data collection. Utilizing the keywords of 535 trending topics on the Chinese Internet between November 25th, 2024 and January 2nd, 2025, we crawl site data of relevant short-videos with durations of less than 5 minutes from the 5 biggest Chinese platforms, i.e., Douyin, Kuaishou, Xigua, Toutiao, and Bilibili. Furthermore, we trace the interactive information of all videos until more than two weeks after their post. Since almost all short-videos peak in interactive indicators and lose popularity within two weeks, our collected data is enough for assessing the long-term propagation influence of short-videos (as further analyzed in Appendix A.4).
3.1. Cross-platform Propagation Influence Annotation
@@
Finally, we design a stepped classification criteria to rate all short-videos in XS-Video into influence levels 0 to 9, using the aligned multi-dimensional indicators . The criteria are designed to make sure that the influence levels follow a natural distribution (Rashid et al., 2005; Cui et al., 2011; Liu et al., 2016). The details of our classification criteria are explained in Appendix A.2.
@@
(d) Propagation influence levels of samples
3.2. Data Analysis
@@
More information about our dataset is provided in Appendix A.
3.3. Short-video Propagation Graph
Utilizing the entities and relations present in XS-Video, we construct a huge propagation graph with the label set . Here is the vertex set, containing 5 million nodes of videos, platforms, topics, titles, descriptions, comments, various interactive indicators, etc. is the edge set, containing 1.7 billion directed edges between nodes, which represent the belonging relations between videos and their attributions, common author relations between videos, common topic relations between videos, chronological relations between samples of the same videos, etc. Specially, is the set of all video nodes and is the set of annotated propagation influence levels of all videos. We analyze the constructed huge propagation graph in detail in Appendix A.3 and Tables 5-6.
4. Method
Short-video propagation network is typical graph-structured data, where edges (such as video-author and video-video) represent complex propagation relations between entries. State-of-the-art methods for analyzing graph-structured data usually adopt Graph Neural Networks (GNNs) (Hong et al., 2020; Yussif et al., 2023; Hu et al., 2024a). Since Large Language Models (LLMs) have recently shown powerful reasoning ability for textual data, researchers have attempted to adapt LLMs to understand graph-structured data (Tang et al., 2024; Li et al., 2024a; Wang et al., 2024b). However, existing research focuses on textual graphs where the information of nodes and edges can be transformed into text, thereby directly leveraging the textual reasoning ability of LLMs. These methods cannot be directly applied to our heterogeneous propagation graph of short-videos. To analyze the short-video propagation graph, where nodes can be texts, times, scalars, and videos, we propose a novel Large Graph Model (LGM) named NetGPT NetGPT fuses open-source LLM (i.e., Qwen2-VL (Wang et al., 2024a) in our experiments) and heterogeneous GNN to effectively comprehend and analyze the short-video propagation graph. Figure 5 illustrates the framework of our three-stage method.
4.1. Heterogeneous Graph Pretraining
Given the short-video propagation graph formulated in Section 3.3 and Tables 5-6, we first extract the raw features of all heterogenous nodes in . For each node of the video modality, we utilize the ViT (Dosovitskiy, 2020) in Qwen2-VL-7B-Instruct444https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct to extract the video feature with an average pooling operation, as follows:
where is the average pooling operation that computes the average feature of the spatial-temporal features outputted by the ViT. For each node of the text modality (including platforms, topics, titles, and descriptions), we utilize the RoBERTa (Liu et al., 2019; Cui et al., 2021) of the Chinese-RoBERTa-wwm-ext-large version555https://huggingface.co/hfl/chinese-roberta-wwm-ext-large to extract the text feature, as follows:
For each node of the time modality (including times and ctimes), as is formulated in the timestamp format, we utilize Sinusoidal Encoding (Vaswani et al., 2017) to compute the time feature, as follows:
where is the output feature of the node , is the dimension of the feature, and is a 2-layer RGCN (Schlichtkrull et al., 2018). Then, we add a 1-layer predictor layer and pretrain the GNN with the predictor by Smooth L1 loss (Girshick, 2015), as follows:
4.2. Supervised Language Fine-tuning
where and are trainable parameters, and is the token embedding dimension of the LLM. Next, we construct the input instruction for LLM, which is “You are a helpful assistant. Graph: . Please read the following json format information about the short-video and predict the final propagation influence level (levels 0 9) of the short video: ”, where is the textual information of the sample (as shown in Figure 2 except the influence level) in the json format. More details of our instruction are introduced in Appendix B.1. Then, we obtain the input token embeddings of , where the embedding of is computed by Equations 7 and 9, and the embeddings of other textual tokens are obtained by the embedding layer of the LLM. The ground truth output of this stage is constructed as response =“The influence level of this short-video is ”, where is the ground truth influence level of . The optimization objective of this stage is maximizing the generation probability of with the input , as follows:
where is the generation probability function of the LLM. Following the original implementation of Qwen2-VL (Wang et al., 2024a), we leverage the cross-entropy loss to optimize the Equation 10. Specially, to preserve the pretrained knowledge and avoid the model corruption introduced by fusing GNN and LLM, we only train the parameters and of the graph feature projector, while keeping all parameters of the pretrained GNN and LLM frozen.
4.3. Task-oriented Predictor Fine-tuning
5. Experiments
More experimental details are included in Appendix C.
5.1. Experimental Setup
Baseline methods. We adopt 10 state-of-the-art baseline methods, including 4 GNNs, 4 LLMs, and 2 multimodal LLMs. The adopted 4 GNN baselines are GCN (Kipf and Welling, 2017), HAN (Wang et al., 2019), HetSANN (Hong et al., 2020), and RGCN (Zhang et al., 2024b). The adopted 4 LLM baselines are Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), InternLM2.5-7B (Cai et al., 2024), Llama-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024). We further adopt 2 multimodal LLM baselines that can comprehend the video modality, i.e., Llava-Next-Video-7B (Li et al., 2024b) and Qwen2-VL-7B-Instruct (Wang et al., 2024a). To conduct a fair comparison, we use the same input graph, predictor module, and optimization loss introduced in Section 4.1 for all GNN baselines. For LLM and multimodal LLM baselines, we use the same input instructions, output response, and optimization loss introduced in Section 4.2, where the information of the short-video is organized in the json format. Additionally, we input the video data into multimodal LLM baselines.
5.2. Metrics
Accuracy (ACC). Following widely-used evaluation for the classification task (Haralick et al., 1973; Krizhevsky et al., 2012; Qian et al., 2023), we compute the Accuracy (ACC) of the predicted results as follows:
where is the rounding function.
Mean Squared Error (MSE). Following widely-used evaluation for the regression task (Draper, 1998; Krizhevsky et al., 2012; Tang et al., 2022), we compute the Mean Squared Error (MSE) of the predicted results as follows:
Model | Input | ACC | MSE | MAE |
GCN | Graph | 0.4474 | 1.0623 | 0.7599 |
HAN | Graph | 0.2619 | 2.6666 | 1.2724 |
HetSANN | Graph | 0.5078 | 0.8917 | 0.6803 |
RGCN | Graph | 0.6313 | 0.7801 | 0.5844 |
Mistral-v0.3 | Text | 0.5387 | 2.1000 | 0.8123 |
InternLM2.5 | Text | 0.5268 | 2.1110 | 0.8064 |
Llama-3.1 | Text | 0.5290 | 2.1215 | 0.8081 |
Qwen2.5 | Text | 0.5469 | 2.0820 | 0.7688 |
Llava-Next-Video | Text+Video | 0.5694 | 1.8503 | 0.7315 |
Qwen2-VL | Text+Video | 0.5884 | 1.6820 | 0.6629 |
NetGPT (Ours) | Graph+Text | 0.6777 | 0.7169 | 0.5457 |
5.3. Results and Discussions
The experimental results on our XS-Video dataset are shown in Table 2. From the results, we have the following observations:
Model | ACC | MSE | MAE |
NetGPT-V | 0.6564 | 0.7415 | 0.5669 |
NetGPT-VV | 0.5889 | 0.8935 | 0.6386 |
NetGPT-IV | 0.2805 | 2.2839 | 1.1688 |
NetGPT-CV | 0.6464 | 0.8249 | 0.5868 |
NetGPT-SLF | 0.6552 | 0.7366 | 0.5605 |
NetGPT-2B | 0.6695 | 0.7307 | 0.5563 |
NetGPT (Ours) | 0.6777 | 0.7169 | 0.5457 |
(a) Long-term prediction (observation periods days)
(b) Median-term prediction (observation periods days)
(c) Short-term prediction (observation periods days)
5.4. Ablation Study
As shown in Table 3, we have the following observations: (1) NetGPT-V performs worse than NetGPT, which shows the importance of considering the video content information. (2) NetGPT-VV, NetGPT-IV, and NetGPT-CV perform worse than NetGPT, which shows the edges from video nodes, interactive information nodes, and comment content nodes to video nodes are significant. (3) NetGPT-SLF underperforms NetGPT, which verifies the effectiveness of our three-stage training mechanism. (4) With a smaller LLM backbone, NetGPT-2B performs worse than NetGPT, which indicates the effectiveness of leveraging pretrained knowledge of larger LLMs.
5.5. Results of Short- and Long-term Prediction
(a) ACC in training procedure
(b) MSE in training procedure
5.6. Convergence on Different Platforms
6. Discussion and Conclusion
Appendix A More Information about Dataset
In this section, we provide more information on our XS-Video dataset.
A.1. Data Collection and Cleaning
A.2. Criteria for Propagation Influence Level Annotation
To evaluate the propagation influence of short videos, our research adopts the cross-platform metrics alignment method described in 3.1 to systematically align the multidimensional interaction metrics accumulated over a two-week period. As shown in Table 4, the criteria for dividing the propagation influence level are as follows: when all interactive indicators of a short-video are 0, its influence level is defined as level 0; for the division of the influence level from 1 to 9, a short video is determined to belong to this influence level when its interactive indicator on any of the five dimensions are more than the corresponding level thresholds. The criteria are designed to make sure that the influence levels follow a natural distribution (Rashid et al., 2005; Cui et al., 2011; Liu et al., 2016), as shown in Figure 4(d).
A.3. Short-video Propagation Graph
Utilizing the entities and relations present in XS-Video, we construct a huge propagation graph consisting of 5.5 million nodes of videos, topics, comments, various interactions, etc., and 1.7 billion directed edges between these nodes. The information about nodes and edges in our graph is shown in Tables 5 and 6, respectively. The short-video propagation graph is a typical heterogenous graph with both heterogenous nodes and edges. In brief, the nodes include the video nodes, the static attribution nodes of videos, and the dynamic interactive nodes of videos. The edges include the edges from the attribution nodes to videos, the interactive nodes to videos, and the video nodes to video nodes.
Views | Likes | Shares | Collects | Comments | |
Level 0 | 0 | 0 | 0 | 0 | 0 |
Level 1 | ¿0 | ¿0 | ¿0 | ¿0 | ¿0 |
Level 2 | ¿100 | ¿10 | ¿10 | ¿10 | ¿10 |
Level 3 | ¿100000 | ¿20 | ¿20 | ¿20 | ¿20 |
Level 4 | ¿1000000 | ¿80 | ¿80 | ¿80 | ¿80 |
Level 5 | ¿1000000 | ¿300 | ¿300 | ¿300 | ¿300 |
Level 6 | ¿1000000 | ¿1000 | ¿1000 | ¿1000 | ¿1000 |
Level 7 | ¿10000000 | ¿3000 | ¿3000 | ¿3000 | ¿3000 |
Level 8 | ¿100000000 | ¿10000 | ¿10000 | ¿10000 | ¿10000 |
Level 9 | ¿200000000 | ¿50000 | ¿50000 | ¿50000 | ¿50000 |
(a) Average increase percentage of likes
(b) Number of tokens
A.4. Statistics of Data

### v2

@@
Abstract
Short-video platforms have gained immense popularity, captivating the interest of millions, if not billions, of users globally. Recently, researchers have highlighted the significance of analyzing the propagation of short-videos, which typically involves discovering commercial values, public opinions, user behaviors, etc. This paper proposes a new Short-video Propagation Influence Rating (SPIR) task and aims to promote SPIR from both the dataset and method perspectives. First, we propose a new Cross-platform Short-Video (XS-Video) dataset, which aims to provide a large-scale and real-world short-video propagation network across various platforms to facilitate the research on short-video propagation. Our XS-Video dataset includes 117,720 videos, 381,926 samples, and 535 topics across 5 biggest Chinese platforms, annotated with the propagation influence from level 0 to 9. To the best of our knowledge, this is the first large-scale short-video dataset that contains cross-platform data or provides all of the views, likes, shares, collects, fans, comments, and comment content. Second, we propose a Large Graph Model (LGM) named NetGPT, based on a novel three-stage training mechanism, to bridge heterogeneous graph-structured data with the powerful reasoning ability and knowledge of Large Language Models (LLMs). Our NetGPT can comprehend and analyze the short-video propagation graph, enabling it to predict the long-term propagation influence of short-videos. Comprehensive experimental results evaluated by both classification and regression metrics on our XS-Video dataset indicate the superiority of our method for SPIR. Our dataset and code will be open upon acceptance.
Index Terms:
Short-video propagation, dataset, large graph model1 Introduction
Short-video platforms, such as Tiktok (aka Douyin)111https://www.tiktok.com/, Kwai (aka Kuaishou)222https://www.kwai.com/, and Xigua333https://www.ixigua.com/, have gained immense popularity, captivating the interest of millions, if not billions, of users globally. The users are capable of publishing their activities, opinions, and thoughts by posting, commenting, and sharing various forms of online content, including text, images, and videos. Consequently, these overwhelming user behaviors and user-generated content have formed a huge and intricate propagation network, harboring a wealth of information about our society. Recently, researchers have highlighted the significance of analyzing the propagation of short-videos [1, 2, 3], which typically involves discovering commercial values, public opinions, user behaviors, etc. Traditional research mainly focuses on forecasting the popularity of videos [4, 5, 6] measured by a single indicator, such as the number of views or likes. However, real-world propagation behaviors are more complex and can include sharing, collecting, commenting, etc.
@@ Abstract
| Daum [7] | 1 | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | 196,037 | 196,037 | Unknown |
| Youtube [8] | 1 | ✓ | ✓ | ✗ | ✗ | ✓ | ✓ | ✗ | 1,761 | 1,761 | Unknown |
| NUS [4] | 1 | ✓ | ✓ | ✗ | ✗ | ✓ | ✓ | ✗ | 303,242 | 303,242 | 98,166 |
| Xigua [5] | 1 | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | 11,219 | 11,219 | 2,664 |
| MicroLens [2] | 1 | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | 19,738 | 19,738 | 100,000 |
@@
To address the above limitation, we attempt to evaluate the propagation influence by simultaneously considering as many interactive indicators as possible, including views, likes, shares, collects, comments, etc. Moreover, as shown in Figure 1, we propose a new task named Short-video Propagation Influence Rating (SPIR), which aims to predict the long-term propagation influence of a newly posted short-video, under the background of a huge propagation network. Different from traditional popularity prediction that forecasts a single interactive indicator over a short period (e.g., several hours [6, 1]), our SPIR predicts the comprehensive propagation influence assessed through multi-dimensional interactions accumulated over a long period. SPIR requires a model to thoroughly analyze the multimodal content features, interactive information, and propagation relationships within a short-video. Subsequently, a rating from 0 to 9 is assigned to represent the propagation influence of this video that can be achieved over a long period. The development of SPIR systems can potentially benefit a wide spectrum of real-world applications on short-video platforms, including advertising [9, 10], recommendation [11, 12], and democracy [13, 14]. In this paper, we attempt to promote the research of SPIR from both the data and model perspectives, proposing the first cross-modal short-video propagation dataset and a novel large graph model for SPIR.
@@
To promote the research of SPIR and short-video propagation analysis, we propose a large-scale Cross-platform Short-Video (XS-Video) dataset, where videos are collected from the 5 biggest Chinese platforms (i.e., Douyin, Kuaishou, Xigua, Toutiao, and Bilibili). Though hot topics and content are typically discussed and replicated across platforms, almost all existing datasets collect data from a single platform and lack a broad scope of short-videos. As compared in Table I, our XS-Video dataset significantly surpasses the existing datasets in terms of platform numbers and interactive indicator coverage, while also containing comment content and more samples and users. Specifically, we collect 381,926 states/samples of 117,720 short-videos within several days after their publication, belonging to 535 topics. Different states/samples of the same video are distinguished by their sampling times, as well as the interactions and comment content at those times. Furthermore, we annotate the propagation influence level of all short-videos according to their multi-dimensional interactive indicators over a two-week period following their publication, since almost all short-videos peak in interactions within two weeks. We demonstrate a short-video sample in the XS-Video dataset in Figure 2. Moreover, utilizing the entities and relations present in XS-Video, we construct a huge propagation graph consisting of 5.5 million nodes of videos, topics, comments, various interactions, etc., and 1.7 billion directed edges between these nodes. More details about our dataset and the constructed propagation graph are provided in Section 3.6. Compared to existing datasets, XS-Video provides more comprehensive information on video content, post, interactions, and comments, which is essential for the research on short-video propagation, including our SPIR task.
@@
We build a comprehensive benchmark for SPIR by evaluating state-of-the-art methods, including Graph Neural Networks (GNNs), Large Language Models (LLMs), and multimodal LLMs. Surprisingly, these state-of-the-art methods are far from satisfactory for short-video propagation analysis. To improve the accuracy of SPIR and leverage the advancements in LLMs, we propose a Large Graph Model (LGM) named NetGPT. Our model can capture both the propagation network structure and heterogenous short-video features while utilizing the pretraining knowledge to predict the propagation influence level. LLMs have recently shown powerful reasoning ability for text, images, and even videos [15, 16, 17, 18, 19]. Current research has explored adapting LLMs for analyzing textual graph data [20, 21, 22]. However, bridging heterogeneous graph data and LLMs is still an open problem, especially as we cannot input our huge propagation graph into an LLM. Our experiments also reveal that LLMs cannot directly learn the graph-structured relations in short-video propagation data, which results in poor performance for SPIR. To address this limitation, we propose a three-stage training mechanism for establishing an effective LGM, including Heterogeneous Graph Pretraining, Supervised Language Fine-tuning, and Task-oriented Predictor Fine-tuning. Finally, we can empower LLMs (e.g., Qwen2-VL [18] in our experiments) to comprehend and analyze the short-video propagation graph. Our NetGPT model significantly outperforms state-of-the-art baselines, including Graph Neural Networks (GNNs), LLMs, and multimodal LLMs on our XS-Video dataset.
@@
We propose the Short-video Propagation Influence Rating (SPIR) task, which aims to predict the long-term propagation influence of a newly posted short-video in a huge propagation network. Different from traditional popularity prediction that forecasts a single interactive indicator over a short period (e.g., several hours [6, 1]), SPIR predicts the comprehensive propagation influence assessed through multi-dimensional interactions over a long period.
@@
2 Related Work
2.1 Short-Video Propagation Influence Analysis
Existing research on short-video propagation influence analysis mainly focuses on the popularity prediction of short-videos, such as predicting the views of newly posted videos [23, 24, 25, 26, 27]. For example, Chen et al. [4] focus on the short-video representation and present a transductive multi-modal learning model for short-video popularity prediction. Xie et al. [5] propose a multimodal variational encoder-decoder framework that considers the uncertain factors as the randomness for the mapping from the multimodal features to the popularity. Xie et al. [6] propose a hierarchical multimodal variational encoder-decoder to predict the popularity of short-videos by comprehensively leveraging user information and the short-video content in a hierarchical fashion. Zhou et al. [2] present a multimodal retrieval-augmented popularity prediction model that improves prediction accuracy using relevant retrieved information.
@@ 2.1 Short-Video Propagation Influence Analysis
2.2 Large Graph Model
Recent advancements in Large Language Models (LLMs) have showcased their impressive reasoning abilities not only with text but also with images and videos [28, 29, 30, 31]. This progress has led to an increasing focus on adapting LLMs for analyzing graph-structured data, especially textual graphs. For instance, Li et al. [21] utilize LLM prompting on textual graphs, offering substantial potential to enhance graph transfer capabilities across diverse tasks and domains. Tang et al. [20] integrate LLMs with graph structural knowledge through graph instruction tuning, which includes a text-graph grounding component, a dual-stage instruction tuning approach, and a graph-text alignment projector. Wang et al. [22] propose Build-a-Graph Prompting and Algorithmic Prompting, two instruction-based approaches to enhance LLMs in solving natural language graph problems.
@@ 2.2 Large Graph Model
3 Dataset
Analyzing propagation patterns on short-video platforms can aid in discovering high-value content, public opinions, and user preferences, thereby creating commercial and social value. Our proposed dataset aims to provide a realistic and large-scale benchmark for the research on short-video propagation. As briefly shown in Table I and Figure 2, compared to previous short-video propagation datasets [7, 8, 4, 5, 2], our dataset has three key advantages: (1) While existing datasets separately collect data from a single short-video platform, we collect short-videos from the 5 biggest Chinese platforms, which provide cross-platform propagation information; (2) Instead of focusing on a subset of interactive indicators, we gather a wide range of indicators such as views, likes, shares, collects, fans, and comments, which facilitate deep analysis of video interactions; (3) Our dataset contains more samples and users, while providing more comprehensive short-video content, including videos, titles, descriptions, topics, and the related comment content. Figure 3 demonstrates a brief construction procedure of our dataset.
3.1 Data Collection and Cleaning
Utilizing the keywords of 535 trending topics on the Chinese Internet between November 25th, 2024 and January 2nd, 2025, we crawl site data of relevant short-videos with durations of less than 5 minutes from the 5 biggest Chinese platforms, i.e., Douyin, Kuaishou, Xigua, Toutiao, and Bilibili. Furthermore, we trace the interactive information of all videos until more than two weeks after their post. Since almost all short-videos peak in interactive indicators and lose popularity within two weeks, our collected data is enough for assessing the long-term propagation influence of short-videos (as further analyzed in Appendix 3.4).
3.2 Cross-platform Propagation Influence Annotation
@@
Finally, we design a stepped classification criteria to rate all short-videos in XS-Video into influence levels 0 to 9, using the aligned multi-dimensional indicators . The criteria are designed to make sure that the influence levels follow a natural distribution [32, 33, 34]. The details of our classification criteria are explained in Appendix 3.3.
@@
(d) Propagation influence levels
(e) Average increase percentage of likes
(f) Number of tokens
| Views | Likes | Shares | Collects | Comments | |
| Level 0 | 0 | 0 | 0 | 0 | 0 |
| Level 1 | 0 | 0 | 0 | 0 | 0 |
| Level 2 | 100 | 10 | 10 | 10 | 10 |
| Level 3 | 100000 | 20 | 20 | 20 | 20 |
| Level 4 | 1000000 | 80 | 80 | 80 | 80 |
| Level 5 | 1000000 | 300 | 300 | 300 | 300 |
| Level 6 | 1000000 | 1000 | 1000 | 1000 | 1000 |
| Level 7 | 10000000 | 3000 | 3000 | 3000 | 3000 |
| Level 8 | 100000000 | 10000 | 10000 | 10000 | 10000 |
| Level 9 | 200000000 | 50000 | 50000 | 50000 | 50000 |
3.3 Criteria for Propagation Influence Level Annotation
To evaluate the propagation influence of short videos, our research adopts the cross-platform metrics alignment method described in 3.2 to systematically align the multidimensional interaction metrics accumulated over a two-week period. As shown in Table II, the criteria for dividing the propagation influence level are as follows: when all interactive indicators of a short-video are 0, its influence level is defined as level 0; for the division of the influence level from 1 to 9, a short video is determined to belong to this influence level when its interactive indicator on any of the five dimensions are more than the corresponding level thresholds. The criteria are designed to make sure that the influence levels follow a natural distribution [32, 33, 34], as shown in Figure 4(d).
3.4 Data Analysis